# Microscopic Mechanism of Anyon Superconductivity Emerging from Fractional Chern Insulators

### v1

@@ I Introduction
Anyons, quasiparticles with fractional statistics, and often with fractional charge, are the emergent low-energy excitations of strongly correlated systems with topological order [1, 2]. The first experimental realization of anyons occurred in fractional quantum Hall (FQH) states, discovered in fractionally filled Landau levels [3]. However, charged anyons in the Landau levels of conventional FQH states lack kinetic energy, due to the strong magnetic field. This enables the formation of incompressible FQH plateaus, but restricts dynamics and prohibits the emergence of competing phases, such as superconductors. Yet, a dispersion can be imprinted on anyons when a periodic potential is applied to a Landau level [4, 5] or when the fractional state is realized from a Chern band. It was realized early on that a gas of charged anyons with a finite dispersion may form a superconductor [6, 7, 8, 9, 10, 11, 12, 13]. Recent experiments with two-dimensional semiconductor heterostructures have indeed realized both fractional Chern insulators stabilized by magnetic fields [14, 15] and Fractional Quantum Anomalous Hall (FQAH) states at zero fields [16, 17, 18, 19, 20]. FQAH states are thus realized on a lattice and are often stable down to zero magnetic field. These properties endow the anyons with finite kinetic energy, opening up the possibility for the formation of competing states, not available in continuum Landau levels [5, 21, 22, 23, 24]. Several recent developments in moiré materials [25], including the observation of superconductivity in MoTe2 [26], the same material in which FQAH states have been realized [16, 17, 18, 19], has rekindled interest in connections between superconductivity and anyons [24, 23, 27].
@@
Simply put, our mechanism invokes a third phase, in addition to the FCI and superconductor. This phase is an insulator but with semion topological order, which for general filling breaks translation symmetry—hence we call it a semion crystal (SX). We will argue that if the FCI is in proximity to the SX, more precisely if it is near a quantum critical point into that phase, then one is guaranteed to have the correct anyon energetics to favor superconductivity. For this reason, we refer to our approach as “anyon superconductivity from topological criticality.” We note that such a crystalline phase is quite natural in the FCI context. It has previously come up to explain insulators with vanishing Chern response observed in MoTe2 [28] and a closely related insulator appeared in the phase diagram of a Hubbard-Hofstadter model [29], but at integer filling. While not essential, these discussions are most transparent when the FQH states arise from electrons with a two-component spin index, which may correspond to physical spin or an approximate pseudo-spin, such as a layer or valley index [30, 29]. In the following, we refer to this additional flavor simply as spin. At the critical point of the topological transition, the charge gap collapses, while the spin sector remains gapped throughout both the FCI phase and the semion crystal reached by strengthening repulsive interactions. This quantum criticality thereby enforces a hierarchy in which the lowest-energy charge carriers are spin singlets. Near criticality, therefore, anyons with even numerator charge cost less energy, guaranteeing that the cheapest anyons fuse to a spinless boson (eg. a Cooper pair) rather than to a spinful electron. Doping either side of the critical point over a broad parameter window should thus yield charge-2e superconductivity, possibly coexisting with translation-symmetry breaking inherited from the crystal (see Fig. 1).
Concrete realization and main results: We consider realizing spin-singlet Halperin FQH states in flat Chern bands. The Halperin states are labeled by the integers $(mmn)$ [31] where for electrons $m$ is an odd integer. Further, we focus on spin-singlet states that arise in the special case of $m=n\pm 1$, and $n\in 2\mathbb{Z}$ [32, 33, 34]. We discuss these states at general fillings of the Chern band $\nu=\frac{2}{2n\pm 1}$ with a parton-vortex approach, within which we can describe the phases FCI, the semion crystal, the superconductor, and the transitions between them.
@@
We further discuss how the mechanism generalizes to fillings $\nu=2/(2n\pm 1)$. Although $\nu=2/3$ enjoys special simplifications, the same logic of the energetic hierarchy applies: spin singlet, bosonic charge-2e anyons emerge as the preferred low-energy carriers near the corresponding topological transitions. The transition between the FCI and SX can be re-expressed as a plateau transition of Cooper pairs from a hierarchy state $\nu_{\rm Cooper}=1/(\mp 2-n^{-1})$ to $\nu_{\rm Cooper}=\mp 1/2$, providing an intuitive bosonic description of all three phases—FCI, semion crystal, and superconductor—and clarifying their interrelations. Interestingly, the relative sign of the charge Hall conductance in the FCI and the sense of propagation of the chiral edge mode in the semion crystal alternates with the ±$\pm 1$ choice in the fractions above. Doping both the FCI and the semion crystal near the transition favors the formation of superconductors, which in turn inherit the alternating chirality of the semion crystal. At these more general fillings, we find that a generic pathway to charge-2 superconductivity without residual topological order exists —provided the superconductor spontaneously breaks translation symmetry. Translationally invariant charge-2e superconductors can also appear upon doping FCIs, yet within our analysis these phases retain an additional topological sector and should be identified as SC states, with $\nu=2/3$ being the exception.
@@ II Phases and Phase transitions at fixed filling: Fractional Chern Insulator and semion crystal
In this section, we show that favorable conditions for anyon superconductivity emerge in isolated Chern bands that host a (pseudo-)spin flavor, such as real spin, a layer, or valley index. A microscopic mechanism for forming anyon superconductivity is then provided when strong repulsion drives these fractionally filled states near competing topological CDW states. Based on parton constructions, we develop a field-theoretic description of such a transition from a Fractional Chern Insulator to the semion crystal (describing a CDW state atop of which a chiral spin liquid forms). This generalizes previous results on the topological transition between an integer quantum Hall state and a chiral spin liquid [29, 35, 36, 27] to fractional fillings.
@@ II.1 Fractional Chern insulators
We find that the $\mathbf{l}_{\uparrow}=(1,0)^{T}$ and $\mathbf{l}_{\downarrow}=(0,1)^{T}$ anyons have electric charge $-e/(2n\pm 1)$. These anyons with the smallest electric charge carry the same spin $S_{z}=\pm 1/2$ quantum numbers as the original electrons, which is a special feature of the $(n\pm 1,n\pm 1,n)$ states, as it makes their topological order compatible with an $SU(2)$ spin rotation symmetry. On the other hand, $\mathbf{l}_{2}=(1,1)^{T}$ anyons carry vanishing spin $S_{z}=0$ and have twice the charge $Q=-2e/(2n\pm 1)$.
@@
Now, the $a$ are conventional gauge fields, and recall that $n$ is even. Integrating out $a$ enforces the constraint $\beta=\alpha_{\uparrow}+\alpha_{\downarrow}$ and we obtain Eq. (1), with K-matrix Eq. (2) for $m=n\pm 1$. The two different signs are understood to arise from spinons forming Chern bands with $C=\pm 1$. For states of the form $(n-1,n-1,n)$, the chirality of the spinons is opposite to that induced by the background flux, resulting in counter-propagating spin and charge edge modes, which is why they are commonly referred to as $(\overline{n-1},\overline{n-1},n)$ states.
@@ II.2 Semion crystal
$\mathcal{L}_{\text{SX}}=\mp\frac{2}{4\pi}\alpha\wedge\mathrm{d}\alpha+\frac{2}{4\pi}A_{s}\wedge\mathrm{d}\alpha,$ (14)
@@ II.3 Interpretation of the FCI to semion crystal transition as bosonic Cooper pair transition
Here we point out that transitions like (1¯1¯2)semion $(\bar{1}\bar{1}2)\leftrightarrow\text{semion crystal}$ can be interpreted as purely bosonic plateau transitions of Cooper pairs. This immediately suggests that there are Cooper pairs near the transition. A similar argument was made for the integer quantum Hall (with $\sigma_{xy}=2e^{2}/h$) to CSL transition in Ref. [27]. Here, we show that a related picture emerges for all the fractional electronic states we consider in this paper as well.
To illustrate this, consider the transition from the $(\bar{1}\bar{1}2)$ FQH state with $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{2/3,-2\}$ to the semion crystal with $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{0,-2\}$. We can trivialize the spin response by stacking a $\nu_{\text{tot}}=2$ IQH state on top of it, composed of spin up and spin down electrons, each in a Chern band with unit Chern number. Since these are integer quantum Hall layers with no bulk anyons, their only effect is to change the edge physics, and hence bulk properties of the original system are unchanged upon stacking. After incorporating the IQH state, this gives: $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{8/3,0\}$ for the FQH and $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{2,0\}$ for the semion crystal. Note, the spin response has been trivialized, which allows for a description of the transition purely in terms of spinless degrees of freedom (see Appendix G for more details). It is nothing but a plateau transition between a Laughlin state of spin singlet Cooper pairs with $\sigma_{xy}^{\text{Cooper}}=\frac{1}{2}\frac{(2e)^{2}}{h}=2\frac{e^{2}}{h}$ and its hierarchy state $\sigma_{xy}^{\text{Cooper}}=(\frac{1}{2}+\frac{1}{6})\frac{(2e)^{2}}{h}=\frac{2}{3}\frac{(2e)^{2}}{h}=\frac{8}{3}\frac{e^{2}}{h}$.
@@ II.4 Effective theory of the critical point
The background fluxes ×$\nabla\times a_{1,2}$ and ×$\nabla\times a$ can now be adjusted such that the fermions $\{f_{\sigma},g_{1},g_{2}\}$ form IQH states at fillings $\{-2,1,1\}$ for $\nu=2/3$ and $\{2,1,1\}$ for $\nu=2/5$. At the same time, this choice ensures a zero-flux condition for the bosons ×$\nabla\times(A-a)=0$. Once the fermions are in gapped IQH states, they contribute terms of the form
$\mathcal{L}_{f_{\sigma}}[a_{1}]$ =±$=\pm\frac{1}{4\pi}(\alpha_{\sigma}\wedge\mathrm{d}\alpha_{\sigma})+\frac{1}{2\pi}a_{1}\wedge\mathrm{d}\alpha_{\sigma}$ (17)
@@
hence, we find that Eq. (16) takes a simple form. We only quote the result for general $n$444for general fillings $\nu=\frac{2}{2n\pm 1}$ we introduce $n$ additional fermions (recall n is even), see Appendix A.:
@@
Thus, in this extended picture we can now analyze the transition in terms of a standard superfluid to a CDW transition of the bosons in the absence of background flux. For this, it is convenient to go to a dual vortex formulation of $b(x)$ [39, 38, 37]. Writing the boson current as $j_{b}^{\mu}=1/2\pi\epsilon^{\mu\nu\sigma}\partial_{\nu}\beta_{\sigma}$ , boson-vortex duality tells us that the vortex field operator $\tilde{\varphi}(x)$ couples minimally to $\beta$. When bosons are at a nonzero background density, the equations of motion tell us that there is a background flux $\mathrm{d}\beta/2\pi$ equal to the background electron density. This, in turn, leads to $p=2n\pm 1$ valleys for the bosonic vortex field. The Lagrangian describing the phase transition reads:
@@
Charge fluctuations at the critical point are then governed by vortices, or equivalently by spinless anyons. Microscopic details encoded in the vortex interactions $V_{\text{int}}(\tilde{\varphi})$ determine the pattern of vortex field condensation $\tilde{\varphi}_{i}$, allowing for the formation of different types of charge density waves [38]. The special case of the bilayer integer quantum Hall $(110)$ to CSL transition, as discussed in Refs. [29, 35, 36, 27], can be understood from setting $n=0$ within the broader framework developed here. There too superconductivity emerges on doping the topological critical point [27], although all Hall conductances are integral. Here we will see that, surprisingly, an analogous outcome is realized even when the insulators exhibit fractional Hall conductance.
@@ III Finite Doping: Superconductivity
For simplicity we focus on doping of the $(\bar{1}\bar{1}2)$ and $(332)$ FQH state at fillings $\nu=2/3$ and $\nu=2/5$, respectively. In Appendix A, we generalize our arguments to general $(n\pm 1,n\pm 1,n)$ states and show that there is always a pathway to simple charge-2 superconductivity that breaks translation symmetry.
@@ III.1 Doping chargon-vortices into the fractional Chern insulator
which keeps all fermionic partons gapped in integer levels for general $\delta$. Since $\mathrm{d}a/2\pi\rightarrow 0$ as $\delta\rightarrow 0$, the bosons $b$ see excess flux only when the system is doped, ensuring that the density of vortices of the $b$ field is also proportional to $\delta$. The key question is the filling of vortices, i.e. the ratio of their density to the flux they see. This is $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$ and will be derived below, but let us give a simple physical argument for it first. The flux seen by the vortices is just the added charge hence $-\delta$. Now, for the density of vortices, we are working in the low-energy space where the only degrees of freedom are vortices which carry electric charge $32e/$ 3. Thus if we add $\delta$ of charge, this implies a vortex density of $(-\delta)/(-2/3\delta)=3\delta/2$. The ratio of this vortex density to the flux gives us: $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$. We will see that there are natural bosonic quantum Hall states that occur at this filling, which will lead overall to a superconductor. Equations of motion provide the following constraints:
@@
which is just the filling. The first term $-1/3$ gives us three valleys for the vortices into which we can dope. The second term is the flux that must be accommodated by forming a bosonic quantum Hall state. Note, the vortices are at filling: $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$. There are two natural choices for vortices at this filling. At $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$, they could form a Jain state and spontaneously pick a single valley
@@
for the field vector $\left(\alpha_{\uparrow},\,\alpha_{\downarrow},\,\gamma_{1},\,\gamma_{2},\gamma_{3}\right)$ where the first two fields support the 2/3 FQH state, and the last three account for doping. For this theory we have $\det K^{(2/3)}_{\mathrm{doped}}=0$ with zero eigenvalue of eigenvector $(1,1,1,1,1)^{T}$ The K-matrix can be block-diagonalized 888consider $V=\begin{pmatrix}\mathrm{Id}_{4\times 4}&0\\ \vec{1}&1\end{pmatrix}$ (34) with $\det V=1$. with a unimodular matrix $\tilde{K}^{(2/3)}_{\mathrm{doped}}=VK^{(2/3)}_{\mathrm{doped}}V^{T}$ which leaves us with the following theory
@@
where $\alpha$ is the gauge field corresponding to the zero eigenvalue of $K^{3/2}_{\text{doped}}$. Integrating it out leads to the Meissner effect $dA=0$. Thus, Eq. (35) describes a charge-2 superconductor with no residual topological order. Furthermore, the sign of the quantized spin response tells us that this is a ‘d-id’ superconductor, at least at the level of edge modes, with $c_{-}=-2$, i.e. two chiral edge modes propagating opposite to the direction of the charge Hall conductance, consistent with [23].
Finally, let us discuss translation symmetry. This depends on the valley structure of the doped vortices. Forming a singlet of the SU(3) valley symmetry automatically preserves the translational symmetry of the system, since the latter is a subgroup of the former. However, spontaneously choosing a valley and subsequently forming a $p=-3$ Jain state breaks translational symmetry, leading to a SC+CDW phase. Both are possible options that can be realized, which depend on the specifics of the microscopic interaction.
Doping the $\nu=2/5$ state.— To describe the $(332)$ state at $\nu=2/5$, we only need to switch the flux experienced by the spinons, such that they each reside in an IQH state at level $+1$. Keeping all the partons in their respective levels corresponds to the following flux assignment:
@@
When combined with $\mathcal{L}_{\text{FQH}}[a]$ this produces a charge-2 superconductor without residual topological order. As in the case of $2/3$ filling, the superconductor inherits the spin edge modes from the spin sector (or equivalently $c_{-}=+2$), which are now propagating in the same direction as the charge modes of the FQH state. This pattern persists for all singlet FCIs: conventional charge-2 superconductivity can form if it also breaks translations. Appendix A explains this construction in detail and also presents new families of SC states that may arise from the doped FCIs. These translation symmetry-breaking superconductors differ from those proposed in [23] at related fillings.

### v2

@@ I Introduction
Anyons, quasiparticles with fractional statistics, and often with fractional charge, are the emergent low-energy excitations of strongly correlated systems with topological order [83, 69]. The first experimental realization of anyons occurred in fractional quantum Hall (FQH) states, discovered in fractionally filled Landau levels [77]. However, charged anyons in the Landau levels of conventional FQH states lack kinetic energy, due to the strong magnetic field. This enables the formation of incompressible FQH plateaus, but restricts dynamics and prohibits the emergence of competing phases, such as superconductors [74]. Yet, a dispersion can be imprinted on anyons when a periodic potential is applied to a Landau level [33, 75] or when the fractional state results from a Chern band [72, 20, 65, 50, 58]. It was realized early on that a gas of charged anyons with a finite dispersion may form a superconductor [38, 37, 13, 41, 79, 26, 80, 42, 6]. Recent experiments with two-dimensional semiconductor heterostructures have indeed observed both fractional Chern insulators stabilized by magnetic fields [73, 84] and Fractional Quantum Anomalous Hall (FQAH) states at zero fields [5, 91, 53, 85, 46]. FQAH states are thus realized on a lattice and are often stable down to zero magnetic field. These properties endow the anyons with finite kinetic energy, opening up the possibility for the formation of competing states, not available in continuum Landau levels [75, 43, 62, 67, 32]. Several recent developments in moiré materials [22], including the observation of superconductivity in MoTe2 [86], the same material in which FQAH states have been realized [5, 91, 53, 85], has rekindled interest in connections between superconductivity and anyons [32, 67, 9].
@@
Simply put, our mechanism invokes a third phase, in addition to the FCI and superconductor. This phase is an insulator but with semion topological order, which for general filling breaks translation symmetry—hence we call it a semion crystal (SX). We will argue that if the FCI is in proximity to the SX, more precisely if it is near a quantum critical point into that phase, then one is guaranteed to have the correct anyon energetics to favor superconductivity. Such a topological criticality mechanism for superconductivity has recently been discussed for integer Chern states adjacent to chiral spin liquids [9], which provides a modern perspective on anyon superconductivity of charge-e semions [37, 38]. The semion crystal phase is quite natural in the FCI context. It has previously come up to explain insulators with vanishing Chern response observed in MoTe2 [71]. The appearance of a CDW with semion topological order seems surprising at first sight; the following viewpoint may help clarify why it emerges here. In the same triangular lattice Hubbard model with $\pi/2$ flux per plaquette, at unit filling per site, a chiral spin liquid (CSL) is well established [34]. For the case considered in this work, at 2/3 filling the system instead breaks translation symmetry forming a new triangular lattice with a three times larger unit cell, such that each elementary plaquette now encloses $3\pi/2=-\pi/2{\,\rm(mod}\,2\pi\rm)$ flux. The effective lattice again favors a CSL, but with reversed chirality, thus rationalizing both the emergence of the SX and the predicted opposite chirality of the topological order.
While not essential, these discussions are most transparent when the FQH states arise from electrons with a two-component spin index, which may correspond to physical spin or an approximate pseudo-spin, such as a layer or valley index [92, 34]. In the following, we refer to this additional flavor simply as spin. At the critical point of the topological transition, the charge gap collapses, while the spin sector remains gapped throughout both the FCI phase and the semion crystal reached by strengthening repulsive interactions. This quantum criticality thereby enforces a hierarchy in which the lowest-energy charge carriers are spin singlets. Near criticality, therefore, anyons with even numerator charge cost less energy, guaranteeing that the cheapest anyons fuse to a spinless boson (eg. a Cooper pair) rather than to a spinful electron. Doping either side of the critical point over a broad parameter regime should thus yield charge-2e superconductivity, possibly coexisting with translation-symmetry breaking inherited from the crystal (see Fig. 1).
Concrete realization and main results: We consider realizing spin-singlet Halperin FQH states in flat Chern bands. The Halperin states are labeled by the integers $(mmn)$ [21] where for electrons $m$ is an odd integer. Further, we focus on spin-singlet states that arise in the special case of $m=n\pm 1$, and $n\in 2\mathbb{Z}$ [2, 29, 12]. We discuss these states at general fillings of the Chern band $\nu=\frac{2}{2n\pm 1}$ with a parton-vortex approach, within which we can describe the FCI, the semion crystal, the superconductor, and the transitions between them.
@@
We further discuss how the mechanism generalizes to fillings $\nu=2/(2n\pm 1)$. Although $\nu=2/3$ enjoys special simplifications, the same logic of the energetic hierarchy applies: spin singlet, bosonic charge-2e anyons emerge as the preferred low-energy carriers near the corresponding topological transitions. The transition between the FCI and SX can be re-expressed as a plateau transition of Cooper pairs from a hierarchy state $\nu_{\rm Cooper}=1/(\mp 2-n^{-1})$ to $\nu_{\rm Cooper}=\mp 1/2$, providing an intuitive bosonic description of all three phases—FCI, semion crystal, and superconductor—and clarifying their interrelations. Interestingly, the relative sign of the charge Hall conductance in the FCI and the sense of propagation of the chiral edge mode in the semion crystal alternates with the $\pm 1$ choice in the fractions above. Doping both the FCI and the semion crystal near the transition favors the formation of superconductors, which in turn inherit the alternating chirality of the semion crystal. At these more general fillings, we find that a generic pathway to charge-2e superconductivity without residual topological order exists—provided the superconductor spontaneously breaks translation symmetry. Translationally invariant charge-2e superconductors can also appear upon doping FCIs, yet within our analysis these phases retain an additional topological sector and should be identified as SC states, with $\nu=2/3$ being the exception.
@@ II Phases and Phase transitions at fixed filling: Fractional Chern Insulator and semion crystal
In this section, we show that favorable conditions for anyon superconductivity emerge in isolated Chern bands that host a (pseudo-)spin flavor, such as real spin, a layer, or valley index. A microscopic mechanism for forming anyon superconductivity is then provided when strong repulsion drives these fractionally filled states near competing topological CDW states. Based on parton constructions, we develop a field-theoretic description of such a transition from a Fractional Chern Insulator to the semion crystal (describing a CDW state atop of which a chiral spin liquid forms). This generalizes previous results on the topological transition between an integer quantum Hall state and a chiral spin liquid [34, 35, 10, 9] to fractional fillings. While we focus on Halperin states in the following, our arguments can be generalized to the Jain sequence (see Appendix I).
@@ II.1 Fractional Chern insulators
We find that the $\mathbf{l}_{\uparrow}=(1,0)^{T}$ and $\mathbf{l}_{\downarrow}=(0,1)^{T}$ anyons have electric charge $e/(2n\pm 1)$. These anyons with the smallest electric charge carry the same spin $S_{z}=\pm 1/2$ quantum numbers as the original electrons, while $\mathbf{l}_{2}=(1,1)^{T}$ anyons carry vanishing spin $S_{z}=0$ and have twice the charge $Q=-2e/(2n\pm 1)$. Thus, the electron is decomposed of anyons which do not carry a fractionalized spin and the original action of an electronic $\mathrm{SU(2)}$ spin symmetry is preserved on the level of the anyons. This is a special property of the $(n\pm 1,n\pm 1,n)$ states consistent with their spin-singlet nature [2, 49].
@@
Now, the $a$ are conventional gauge fields, and recall that $n$ is even. Integrating out $a$ enforces the constraint $\beta=\alpha_{\uparrow}+\alpha_{\downarrow}$ and we obtain Eq. (1), with K-matrix Eq. (2) for $m=n\pm 1$. The two different signs are understood to arise from spinons forming Chern bands with $C=\pm 1$. For states of the form (n1,n1,n)(n-1,n-1,n), the chirality of the spinons is opposite to that induced by the background flux, resulting in counter-propagating spin and charge edge modes, which is why they are commonly referred to as $(\overline{n-1},\overline{n-1},n)$ states.
@@ II.2 Semion crystal
$\mathcal{L}_{\text{SX}}=\mp\frac{2}{4\pi}\alpha\wedge\mathrm{d}\alpha+\frac{2}{2\pi}A_{s}\wedge\mathrm{d}\alpha,$ (14)
@@ II.3 Interpretation of the FCI to semion crystal transition as bosonic Cooper pair transition
[…]
To illustrate this, consider the transition from the $(\bar{1}\bar{1}2)$ FQH state with $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{2/3,-2\}$ to the semion crystal with $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{0,-2\}$. We can trivialize the spin response by stacking a $\nu_{\text{tot}}=2$ IQH state on top of it, composed of spin up and spin down electrons, each in a Chern band with unit Chern number. Since these are integer quantum Hall layers with no bulk anyons, their only effect is to change the edge physics, and hence bulk properties of the original system are unchanged upon stacking. After incorporating the IQH state, this gives: $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{8/3,0\}$ for the FQH and $\{\sigma_{xy},\sigma_{xy}^{s}\}=\{2,0\}$ for the semion crystal. Note, the spin response has been trivialized, which allows for a description of the transition purely in terms of spinless degrees of freedom (see Appendix H for more details). It is nothing but a plateau transition between a Laughlin state of spin singlet Cooper pairs with $\sigma_{xy}^{\text{Cooper}}=\frac{1}{2}\frac{(2e)^{2}}{h}=2\frac{e^{2}}{h}$ and its hierarchy state $\sigma_{xy}^{\text{Cooper}}=(\frac{1}{2}+\frac{1}{6})\frac{(2e)^{2}}{h}=\frac{2}{3}\frac{(2e)^{2}}{h}=\frac{8}{3}\frac{e^{2}}{h}$.
@@ II.4 Effective theory of the critical point
The background fluxes $\nabla\times a_{1,2}$ and $\nabla\times a$ can now be adjusted such that the fermions $\{f_{\sigma},g_{1},g_{2}\}$ form IQH states at fillings $\{-2,1,1\}$ for $\nu=2/3$ and $\{2,1,1\}$ for $\nu=2/5$. At the same time, this choice ensures a zero-flux condition for the bosons $\nabla\times(A-a)=0$. Once the fermions are in gapped IQH states, they contribute terms of the form
$\mathcal{L}_{f_{\sigma}}[a_{1}]$ $=\pm\frac{1}{4\pi}(\alpha_{\sigma}\wedge\mathrm{d}\alpha_{\sigma})+\frac{1}{2\pi}a_{1}\wedge\mathrm{d}\alpha_{\sigma}$ (17)
@@
hence, we find that Eq. (16) takes a simple form. We only quote the result for general nn444for general fillings $\nu=\frac{2}{2n\pm 1}$ we introduce $n$ additional fermions (recall n is even), see Appendix A.:
@@
Thus, in this extended picture we can now analyze the transition in terms of a standard superfluid to a CDW transition of the bosons in the absence of background flux. As before, the FQH state forms when $b$ condenses. Close to the transition, vortices of the $b$ condensate are natural excitations, which become soft as the transition is approached. Thus, to describe excitations in the FQH state it is convenient to work directly with vortex operators. This is achieved by going to a dual vortex formulation of $b(x)$ [36, 4, 3]. Writing the boson current as $j_{b}^{\mu}=1/2\pi\epsilon^{\mu\nu\sigma}\partial_{\nu}\beta_{\sigma}$ , boson-vortex duality tells us that the vortex field operator $\tilde{\varphi}(x)$ couples minimally to $\beta$. When bosons are at a nonzero background density, the equations of motion tell us that there is a background flux $\mathrm{d}\beta/2\pi$ equal to the background electron density. This, in turn, leads to $p=2n\pm 1$ valleys in the disperison of the bosonic vortex field. The Lagrangian describing the phase transition reads:
@@
[…]
@@ III Finite Doping: Superconductivity
For simplicity we focus on doping of the $(\bar{1}\bar{1}2)$ and $(332)$ FQH state at fillings $\nu=2/3$ and $\nu=2/5$, respectively. In Appendix A, we generalize our arguments to general $(n\pm 1,n\pm 1,n)$ states and show that there is always a pathway to simple charge-2e superconductivity that breaks translation symmetry.
@@ III.1 Doping chargon-vortices into the fractional Chern insulator
which keeps all fermionic partons gapped in integer levels for general $\delta$. Since $\mathrm{d}a/2\pi\rightarrow 0$ as $\delta\rightarrow 0$, the bosons $b$ see excess flux only when the system is doped, ensuring that the density of vortices of the $b$ field is also proportional to $\delta$. The key question is the filling of vortices, i.e. the ratio of their density to the flux they see. This is $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$ and will be derived below, but let us give a simple physical argument for it first. The flux seen by the vortices is just the added charge hence $\delta$. Now, for the density of vortices, we are working in the low-energy space where the only degrees of freedom are vortices which carry electric charge $2e/3$. Thus if we add $\delta$ of charge, this implies a vortex density of $(-\delta)/(-2/3\delta)=3\delta/2$. The ratio of this vortex density to the flux gives us: $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$. We will see that there are natural bosonic quantum Hall states that occur at this filling, which will lead overall to a superconductor. Equations of motion provide the following constraints:
@@
which is just the filling. The first term 1/3-1/3 gives us three valleys for the vortices into which we can dope. The second term is the flux that must be accommodated by forming a bosonic quantum Hall state. Note, the vortices are at filling: $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$. There are two natural choices for vortices at this filling. At $\nu^{\mathrm{total}}_{\mathrm{vortices}}=3/2$, they could form a Jain state and spontaneously pick a single valley
@@
for the field vector $\left(\alpha_{\uparrow},\,\alpha_{\downarrow},\,\gamma_{1},\,\gamma_{2},\gamma_{3}\right)$ where the first two fields support the 2/3 FQH state, and the last three account for doping. For this theory we have $\det K^{(2/3)}_{\mathrm{doped}}=0$ with zero eigenvalue of eigenvector $(1,1,1,1,1)^{T}$ . The K-matrix can be block-diagonalized 888consider $V=\begin{pmatrix}\mathrm{Id}_{4\times 4}&0\\ \vec{1}&1\end{pmatrix}$ (34) with $\det V=1$. with a unimodular matrix $\tilde{K}^{(2/3)}_{\mathrm{doped}}=VK^{(2/3)}_{\mathrm{doped}}V^{T}$ which leaves us with the following theory
@@
where $\alpha$ is the gauge field corresponding to the zero eigenvalue of $K^{3/2}_{\text{doped}}$. Integrating it out leads to the Meissner effect $dA=0$. Thus, Eq. (35) describes a charge-2e superconductor with no residual topological order. Furthermore, the sign of the quantized spin response tells us that this is a ‘d-id’ superconductor, at least at the level of edge modes, with $c_{-}=-2$, i.e. two chiral edge modes propagating opposite to the direction of the charge Hall conductance, consistent with [67].
Finally, let us discuss translation symmetry. This depends on the valley structure of the doped vortices. Forming a singlet of the SU(3) valley symmetry automatically preserves the translational symmetry of the system, since the latter is a subgroup of the former. However, spontaneously choosing a valley and subsequently forming a p=3p=-3 Jain state breaks translational symmetry, leading to a SC+CDW phase. Both are possible options that can be realized, which depend on the specifics of the microscopic interaction.
Doping the $\nu=2/5$ state.— To describe the $(332)$ state at $\nu=2/5$, we only need to switch the flux experienced by the spinons, such that they each reside in an IQH state at level +1+1. Keeping all the partons in their respective levels corresponds to the following flux assignment:
@@
When combined with $\mathcal{L}_{\text{FQH}}[a]$ this produces a charge-2e superconductor without residual topological order. As in the case of $2/3$ filling, the superconductor inherits the spin edge modes from the spin sector (or equivalently $c_{-}=+2$), which are now propagating in the same direction as the charge modes of the FQH state. This pattern persists for all singlet FCIs: conventional charge-2e superconductivity can form if it also breaks translations. Appendix A explains this construction in detail and also presents new families of SC states that may arise from the doped FCIs. These translation symmetry-breaking superconductors differ from those proposed in [67] at related fillings.